# HSIGene: A Foundation Model For Hyperspectral Image Generation

### v1

@@ IV-B2 Quantitative results
We conduct a comprehensive comparison of our model with existing HSI synthesis models, i.e., UBF [27] and UnmixDiff [28]. For fair comparison, the evaluation is performed on 1024 images generated by each model in an unconditional setting. The results of this comparison are presented in Table III, where our approach demonstrates superior performance on most metrics. UBF trains a diffusion model on augmented training datasets which is obtained by performing spectral super-resolution on RGB images. Owing to the inherent differences between RGB images and hyperspectral data, while UBF is capable of generating HSIs, it is challenging to ensure the spectral authenticity of the generated images, resulting in both low sPr and sRec. UnmixDiff, on the other hand, is trained directly on real hyperspectral data. However, its effectiveness is limited by the relatively small size of available hyperspectral datasets and model parameters, restricting its ability to generalize and produce high-quality images across diverse scenarios. In addition, the reliability of the generated spectra is constrained by the performance of the unmixing network in UnmixDiff, which generates abundance maps instead of HSIs directly. Our model addresses these limitations by increasing both model size and training data based on latent diffusion and spatial super-resolution augmentation. Particularly, the spatial super-resolution approach significantly increases the number of data samples while maintaining the spectral reliability of the images. This method not only enhances the diversity of the training data but also ensures that the generated images preserve the spectral characteristics of real hyperspectral data, leading to the best performance among the compared models. In addition, the latent representation generation technique adopted in our model enables a larger model size, leading to improved generation performance under scaling laws.
@@
In addition, comparing with UBF and UnmixDiff, our model supports HSI generation under multiple conditions, which enables more accurate and effective generation. Image generation results with the conditions generated by AID dataset are illustrated in Table IV. As shown in the table, our model demonstrates superior performance when more conditions are provided. For example, when four conditions (i.e., MLSD, HED, sketch and segmentation) are simultaneously applied, our model achieves the best results across metrics including SSIM, CLIPScore, BRISQUE and sPr. The results indicate that additional control inputs allow the model to better guide the generation process, resulting in images that are not only structurally accurate but also of higher fidelity and perceptual quality.
IV-C Augmentation analysis
@@
In this section we provide an analysis of our proposed HSI augmentation method. We verify the effectiveness of our proposed HSI augmentation method in two aspects, including the improvement in HSI generation and the image quality of augmented data.
IV-C1 Improvement in HSI generation
@@
The effectiveness of our data augmentation method is demonstrated through a comparison of models trained with and without data augmentation, as shown in Table V. The results indicate that the model trained with augmented data achieves superior performance in most cases, verifying the effectiveness of our proposed data augmentation strategy. As the number of training samples increases, the model is trained on a more diverse set of conditions, leading to improved generalization performance and the ability to generate images that align with various control conditions. Through the application of diffusion models (i.e., DSRNet), we achieve data augmentation while preserving the high spatial quality of the data. The model trained with augmented data shows slightly lower performance in sPr but higher sRec in some cases. This outcome is intuitive because the resolution of the augmented data is higher than that of the real dataset. As a result, the model trained on augmented data is more likely to generate images with higher resolution than the original HSIs, which can lead to more various spectral profiles comparing to real hyperspectral images. Since spectral evaluation metrics (i.e., sPr and sRec) are calculated between generated and real spectra, and the ground truth spectra of high-resolution images are unknown, the sPr metric of spectral evaluation may be slightly lower than models trained without super-resolved data. In contrast, due to the increased spectral diversity, the sRec metric tends to be higher. Despite this, the model achieves a good balance between sPr and sRec, ensuring the authenticity of the generated spectra. In addition, the model trained with augmented data exhibits better performance in structural similarity and overall perceptual image quality, demonstrating that data augmentation not only enhances the model’s ability to generalize across different conditions but also boosts its overall performance in practical applications.
@@
Figure 12: Visualization comparison of the augmented data which is super-resolved by different approaches.
@@
IV-C2 Image quality of augmented data
To measure the image quality of the augmented data and validate the effectiveness of our super-resolution method, we provide quantitative results of no-reference image quality assessment metrics in Table VI and visual results in Fig. 12. It can be seen that our method significantly outperforms other approaches in image perceptual quality. Our method is able to achieve superior results for several reasons. First, the diffusion model (i.e., DSRNet) adopted in our network could effectively recover image details, which is crucial for high-quality super-resolution. Second, we leverage a large dataset of high-resolution RGB images for training, significantly enhancing the network’s performance by providing a rich source of detail and structure. Moreover, the RGAN network proposed in our work effectively transfers the fine details from the RGB modality into the hyperspectral modality, ensuring that the super-resolved hyperspectral images retain high fidelity and sharpness. Owing to the generative ability of DSRNet, the super-resolved images exhibits richer high-frequency details, leading to more diverse spectral profiles. Therefore, as discussed in previous section, our model performs better in terms of sRec but slightly worse in terms of sPr. Despite this, our model achieves desirable results on both sPr and sRec metrics, confirming the spectral reliability of the augmented data.
IV-D Ablation results
In this section, we discuss the effectiveness of the components of our proposed two-stage super-resolution method. We compare our approach to two alternative situations: directly applying the diffusion model DSRNet proposed in Sec.III-C2 or the RGAN network proposed in Sec.III-C3 for HSI super-resolution and training these networks on real HSI data. For the DSRNet we directly employ the low-resolution HSIs as conditions and for the RGAN the input of RGB-modality is set as zeros. The results are illustrated in Table VII. It can be seen that, DSRNet and RGAN, when used independently, struggle to recover high-frequency details in the images. In contrast, our two-stage method leverages both high-resolution data augmentation and the powerful generative capabilities of the diffusion model. The first stage uses high-resolution RGB images to enhance the training dataset, ensuring that the diffusion model learns to reproduce intricate details of high-resolution RGB bands. In the second stage image details are transferred to HSIs using the RGAN, resulting in superior super-resolution performance. Overall, by combining the strengths of high-resolution data augmentation with the diffusion model’s generative power, our approach effectively recovers high-frequency details and produces the highest quality super-resolved HSIs.
@@ IV-D Ablation results
IV-E Application in downstream tasks
@@
We further validate the effectiveness of our generative model on two downstream tasks: HSI denoising and HSI super-resolution. We augment the training datasets for both tasks using our proposed generative model to improve the performance of existing image restoration models. Test images are degraded with Gaussian noise ($\sigma=0.2$) for denoising and downsampled 4×4$\times$4 × for super-resolution respectively. The Xiongan dataset is partitioned into two parts: one part of size $512\times 512$ for testing, and the remaining part for training. In addition, the image restoration models are also tested on the WHU-Hi-HanChuan dataset [79] which is not used for training our generative model. For training, the Xiongan dataset is cropped into patches whose size is $128\times 128$ pixels with a stride of 32, resulting in approximately 5k training patches. In the augmented data scenario, we used our generative model to synthesize 100 additional HSIs with a resolution of $256\times 256$. These generated images are then also cropped into patches of the size $128\times 128$, producing another 5k training patches. This augmentation effectively enhances the training data, providing a total of 10k patches for training. All training data are degraded in the same setting as test images to generated pair-wise training images. We evaluate the performance using three metrics including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) and Spectral Angle Mapper (SAM). We employ T3SC [80], TRQ3D [81], SERT [59] for denoising evaluation and MCNet [38], SSPSR [40], ESSAformer [41] for super-resolution evaluation. All models are trained for 30 epochs and the batch size is set as 4. The learning rate is set to $10^{-4}$ with AdamW optimizer and decreases to $10^{-5}$ with cosine annealing.
@@
The results are presented in Table VIII and Table IX. As can be seen, when the test set and training set have the same distribution (i.e., the Xiongan dataset), models trained with both augmented and real data exhibit comparable performance to the model trained with only real data. However, when tested on the WHU-Hi-HanChuan dataset, the performance of models trained solely on real data, such as TRQ3D and SERT, deteriorates significantly, while the models trained with both augmented and real data still achieve the best results. Additionally, the model trained solely on augmented data also achieves relatively desirable performance, proving the high quality of our generated data. By augmenting the training data, our approach contributes to more robust and accurate models, demonstrating the broader impact and effectiveness of our work in advancing hyperspectral image processing.
V Conclusion
@@
In this paper, we present a foundation model namely HSIGene for HSI synthesis, which utilizes a latent diffusion model to synthesize HSIs with support for multiple control conditions. To enhance the spatial diversity of the training dataset, we propose a spatial super-resolution based data augmentation method and design a two-stage super-resolution approach to improve the perceptual quality of the augmented images. Extensive experiments demonstrate that our model outperforms existing methods in HSI synthesis and verifies the reliability of the augmented data. The results of two downstream tasks including HSI denoising and HSI super-resolution demonstrate that our model could provide a substantial amount of high-quality data for model training, significantly boosting the performance of downstream tasks.

### v2

@@ 4.2.2 Quantitative results
We comprehensively compare our model with existing HSI synthesis models, i.e., UBF [19] and UnmixDiff [20]. For fairness, the evaluation is performed on 1024 images generated by each model in an unconditional setting. The results of this comparison are presented in Table III, where our approach demonstrates superior performance on most metrics. UBF trains a diffusion model on augmented training datasets which is obtained by performing spectral super-resolution on RGB images. Owing to the inherent differences between RGB images and hyperspectral data, while UBF is capable of generating HSIs, it is challenging to ensure the spectral authenticity of the generated images, resulting in both low sPr and sRec. UnmixDiff, on the other hand, is trained directly on real HSI data. However, its effectiveness is limited by the relatively small size of available hyperspectral datasets and model parameters, restricting its ability to generalize and produce high-quality images across diverse scenarios. In addition, the reliability of the generated spectra is constrained by the performance of the unmixing network in UnmixDiff, which generates abundance maps instead of HSIs directly. Our model addresses these limitations by increasing both model size and training data based on latent diffusion and spatial super-resolution augmentation. Particularly, the spatial super-resolution approach significantly increases the number of data samples while maintaining the spectral reliability of the images. This method not only enhances the diversity of the training data but also ensures that the generated images preserve the spectral characteristics of real hyperspectral data, leading to the best performance among the compared models. In addition, the latent representation generation technique adopted in our model enables a larger model size, leading to improved generation performance under scaling laws.
@@
Besides, compared with UBF [19] and UnmixDiff [20], our model supports HSI generation under multiple conditions, enabling more accurate and effective generation. Image generation results with the conditions generated by the AID dataset are shown in Table IV. As can be seen, our model obtains superior performance when more conditions are provided. For example, when four conditions (i.e., MLSD, HED, sketch and segmentation) are simultaneously applied, our model achieves the best results across metrics including SSIM, CLIPScore, BRISQUE and sPr. The results indicate that additional control inputs allow the model to better guide the generation process, resulting in images that are not only structurally accurate but also of higher fidelity and perceptual quality.
4.3 Augmentation analysis
@@
In this section, we provide an analysis of our proposed HSI augmentation method. We verify the effectiveness of our proposed HSI augmentation method in two aspects, including the improvement in HSI generation and the image quality of augmented data.
4.3.1 Improvement in HSI generation
@@
The effectiveness of our data augmentation method is demonstrated through a comparison of models trained with and without data augmentation, as shown in Table V. The results indicate that the model trained with augmented data achieves superior performance in most cases, verifying the effectiveness of our proposed data augmentation strategy. As the number of training samples increases, the model is trained on a more diverse set of conditions, leading to improved generalization performance and the ability to generate images that align with various control conditions. Through the application of diffusion models (i.e., DSRNet), we achieve data augmentation while preserving the high spatial quality of the data. The model trained with augmented data shows slightly lower performance in sPr but higher sRec in some cases. This outcome is intuitive because the resolution of the augmented data is higher than that of the real dataset. As a result, the model trained on augmented data is more likely to generate images with higher resolution than the original HSIs, which can lead to more various spectral profiles comparing to real hyperspectral images. Since spectral evaluation metrics (i.e., sPr and sRec) are calculated between generated and real spectra, and the ground truth spectra of high-resolution images are unknown, the sPr metric of spectral evaluation may be slightly lower than models trained without super-resolved data. In contrast, due to the increased spectral diversity, the sRec metric tends to be higher. Despite this, the model achieves a good balance between sPr and sRec, ensuring the authenticity of the generated spectra. In addition, the model trained with augmented data exhibits better performance in both structural similarity and overall perceptual image quality, demonstrating that data augmentation not only enhances the model’s ability to generalize across different conditions but also enhances its robustness and effectiveness in practical applications.
@@
Figure 12: Visualization comparison of the augmented data super-resolved by different approaches.
@@
4.3.2 Image quality of augmented data
To measure the image quality of the augmented data and validate the effectiveness of our super-resolution method, we provide quantitative results of no-reference image quality assessment metrics in Table VI and visual results in Fig. 12. It can be seen that our method significantly outperforms other approaches in image perceptual quality. Our method is able to achieve superior results for several reasons. First, the diffusion model (i.e., DSRNet) adopted in our network could effectively recover image details, which is crucial for high-quality super-resolution. Second, we leverage a large dataset of high-resolution RGB images for training, significantly enhancing the network’s performance by providing a rich source of detail and structure. Moreover, the RGAN network proposed in our work effectively transfers the fine details from the RGB modality into the hyperspectral modality, ensuring that the super-resolved hyperspectral images retain high fidelity and sharpness. Owing to the generative ability of DSRNet, the super-resolved images exhibit richer high-frequency details, leading to more diverse spectral profiles. Therefore, as discussed in the previous section, our model performs better in terms of sRec but slightly worse in terms of sPr. Despite this, our model achieves desirable results on both sPr and sRec metrics, confirming the spectral reliability of the augmented data.
4.4 Ablation results
In this section, we discuss the effectiveness of the components of our proposed two-stage super-resolution method. We compare our approach to two alternative situations: directly applying the diffusion model DSRNet proposed in Sec.3.3.2 or the RGAN network proposed in Sec.3.3.3 for HSI super-resolution and training these networks on real HSI data. For the DSRNet we directly employ the low-resolution HSIs as conditions and for the RGAN the input of RGB-modality is set as zeros. The results are illustrated in Table VII. It can be seen that, DSRNet and RGAN, when used independently, struggle to recover high-frequency details in the images. In contrast, our two-stage method leverages both high-resolution data augmentation and the powerful generative capabilities of the diffusion model. The first stage uses high-resolution RGB images to enhance the training dataset, ensuring that the diffusion model learns to reproduce intricate details of high-resolution RGB bands. In the second stage, image details are transferred to HSIs using the RGAN, resulting in superior super-resolution performance. Overall, by combining the strengths of high-resolution data augmentation with the diffusion model’s generative power, our approach effectively recovers high-frequency details and produces the highest quality super-resolved HSIs.
@@ 4.4 Ablation results
4.5 Application in downstream tasks
@@
We further validate the effectiveness of our generative model on two downstream tasks: HSI denoising and HSI super-resolution. We augment the training datasets for both tasks using our proposed generative model to improve the performance of existing image restoration models. Test images are degraded with Gaussian noise ($\sigma=0.2$) for denoising and downsampled 4×4$\times$4 × for super-resolution respectively. The Xiongan dataset is partitioned into two parts: one part of size $512\times 512$ for testing, and the remaining part for training. In addition, the image restoration models are also tested on the WHU-Hi-HanChuan dataset [69] which is not used for training our generative model. For training, the Xiongan dataset is cropped into patches whose size is $128\times 128$ pixels with a stride of 32, resulting in approximately 5k training patches. In the augmented data scenario, we used our generative model to synthesize 100 additional HSIs with a resolution of $256\times 256$. These generated images are then also cropped into patches of the size $128\times 128$, producing another 5k training patches. This augmentation effectively enhances the training data, providing a total of 10k patches for training. All training data are degraded in the same setting as test images to generate pair-wise training images. We evaluate the performance using three metrics including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) and Spectral Angle Mapper (SAM). We employ T3SC [67], TRQ3D [68], SERT [47] for denoising evaluation and MCNet [26], SSPSR [28], ESSAformer [29] for super-resolution evaluation. All models are trained for 30 epochs and the batch size is 4. The learning rate is $10^{-4}$ with AdamW optimizer and decreases to $10^{-5}$ with a strategy of cosine annealing.
@@
The results are presented in Table VIII and Table IX. As can be seen, when the test set and training set have the same distribution (i.e., the Xiongan dataset), models trained with both augmented and real data exhibit comparable performance to the model trained with only real data. However, when tested on the WHU-Hi-HanChuan dataset, the performance of models trained solely on real data, such as TRQ3D and SERT, deteriorates significantly, while the models trained with both augmented and real data still achieve the best results. Additionally, the model trained solely on augmented data also achieves relatively desirable performance, proving the high quality of our generated data. By augmenting the training data, our approach contributes to more robust and accurate models, demonstrating the broader impact and effectiveness of our work in advancing more HSI processing tasks.
5 Conclusion
@@
In this paper, we present a foundation model namely HSIGene for HSI synthesis, which utilizes a latent diffusion model to synthesize HSIs with support for multiple control conditions. To enhance the spatial diversity of the training dataset, we propose a spatial super-resolution based data augmentation method and design a two-stage super-resolution approach to improve the perceptual quality of the augmented images. Extensive experiments demonstrate that our model outperforms existing methods in HSI synthesis and verifies the reliability of the augmented data. The results of two downstream tasks including HSI denoising and HSI super-resolution demonstrate that our model could provide a substantial amount of high-quality data for model training, boosting the performance of downstream tasks.